OpenStudy (anonymous):

What is the solution to the following system of equations? y = -x2 - 5x - 4 y = -x2 + 9x - 18

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

You can try factoring it

OpenStudy (anonymous):

but first get rid of the negative in the front

OpenStudy (anonymous):

so is it y=x2+5x+4 y=x2-9x+18?

OpenStudy (anonymous):

y=(x+1)(x+4) y=(x-6)(x-3)